Lars:

What is the volume of the Sphere? 𝝅𝒌𝒎3 This is the last question I have today, and I'm confused on how I solve this with the formula, would it be 21^3 times 4 / 3?

Lars:

1 attachment
KyledaGreat:

I'll give you a chance to guess the answer to the question about the volume of a sphere with a diameter of 21 cm. What do you think the answer is?

KyledaGreat:

Based on the information provided, can you take a guess at what the volume of the sphere might be? Remember, the formula for the volume of a sphere is V = (4/3)πr^3, where r is the radius of the sphere.

Lars:

@kyledagreat wrote:
Based on the information provided, can you take a guess at what the volume of the sphere might be? Remember, the formula for the volume of a sphere is V = (4/3)πr^3, where r is the radius of the sphere.
12348𝝅𝒌𝒎3 ?

Shadow:

@lars wrote:
@kyledagreat wrote:
Based on the information provided, can you take a guess at what the volume of the sphere might be? Remember, the formula for the volume of a sphere is V = (4/3)πr^3, where r is the radius of the sphere.
12348𝝅𝒌𝒎3 ?
I got the same thing. Just make sure its km^3

Lars:

@shadow wrote:
@lars wrote:
@kyledagreat wrote:
Based on the information provided, can you take a guess at what the volume of the sphere might be? Remember, the formula for the volume of a sphere is V = (4/3)πr^3, where r is the radius of the sphere.
12348𝝅𝒌𝒎3 ?
I got the same thing. Just make sure its km^3
What do you mean by make sure its km^3? Does that mean change the answer to 12348𝝅km^3 ?

Shadow:

\[12345 \pi km^3\] It's just how it's supposed to look. No biggie.

Lars:

@shadow wrote:
\[12345 \pi km^3\] It's just how it's supposed to look. No biggie.
Oh okay. 12345πkm3 is how I should put the answer instead of 12348𝝅𝒌𝒎3?

KyledaGreat:

I apologize for any confusion. In the previous question, I provided the answer in cubic kilometers (km^3), which is a unit of volume. If you were asked for the answer in cubic kilometers, the answer, which is approximately 4850.32 km^3. If you were asked for the answer in cubic meters (m^3), then you can convert the answer to m^3 by multiplying by 1 billion (1 km^3 = 1 billion m^3). So, the answer in cubic meters would be approximately 4.85 × 10^12 m^3. In any case, it is important to include the correct units of measurement when providing an answer to a question.

KyledaGreat:

@lars wrote:
@shadow wrote:
\[12345 \pi km^3\] It's just how it's supposed to look. No biggie.
Oh okay. 12345πkm3 is how I should put the answer instead of 12348𝝅𝒌𝒎3?
Yes, if you were asked to provide the answer in cubic kilometers, then 12345πkm^3 would be the correct format for the answer. Just to clarify, the correct volume of the sphere is approximately 4850.32 km^3, which is calculated using the formula V = (4/3)πr^3, where r is the radius of the sphere. I apologize for any confusion earlier, and I hope this clears things up for you.

Lars:

Ohh okay i think that makes sense. So the final answer is that the sphere's volume is 4850.32 km^3 then and not what i said above?

Lars:

Now I'm even more confused.

Shadow:

All your questions have had _____𝝅𝒌𝒎3

Shadow:

So just don't include pi in the calculation and leave it next to the number. The true volume is 4850.32 km^3 as Kyle said. but if they want the answer in 𝝅𝒌𝒎^3 then you would use the number that doesn't have pi multiplied into it.

Shadow:

It's like if I were to ask for a dish of french fries but I want the ketchup on the side, and not drizzled over it.

Lars:

Ohh okay I see now, because the question isn't asking in cubic kilometers

Shadow:

Well, it's asking for it in 𝝅𝒌𝒎^3, which means they want pi on the side, and not calculated into the answer.

Lars:

Ohh okay then, thank you both